# The TeraZ Mirage: New Physics Lost in Blind Directions

### v1

@@ II Effective Field Theory Fit: bottom up
We emphasize that this is not a complete list of all blind directions, but just the simplest ones we have found with the stated assumptions. As we will see in the next section, this is sufficient to illustrate the blindness of TeraZ to some classes of BSM models, since there are several extensions of the SM that generate each of them.
We further observe that finite one-loop SMEFT effects [24] slightly modify, but do not eliminate, the blind directions identified. In the following section, we systematically examine extensions of the SM that, when matched at tree level onto the SMEFT, give rise to effective operators aligned with these blind subspaces.
@@ III UV extensions of the SMEFT
The complete classification of SM extensions that generate dimension-six SMEFT operators at tree level was presented in Ref. [38], building on earlier work [39, 40, 28]. For clarity, we reproduce in Table 1 only the subset relevant to our analysis, along with the corresponding effective operators they induce. We adopt the coupling notation used in Ref. [38], and note that when multiple heavy fields are present, their contributions to the SMEFT Lagrangian simply add at leading order.
We assume that each heavy field $X$ has a single coupling to the SM, schematically of the form $X\psi_{1}\psi_{2}$, where $\psi_{1}$ and $\psi_{2}$ are SM fermions. The contribution of such a field $X$ to the SMEFT Lagrangian at tree level and dimension six is a linear combination of one or two Warsaw-basis operators with the field content $(\psi_{1})^{2}(\psi_{2})^{2}$. The only pairs of operators that may be generated are $\{\mathcal{O}_{x}^{(1)},\mathcal{O}_{x}^{(8)}\}$, with $x=qu$, $qd$ and $ud$. The effects of the $\mathcal{O}_{x}^{(8)}$ operators on EWPO are all negligible. Thus, each heavy field we consider effectively generates a single operator relevant for EWPO.
Our aim is to identify which tree-level completions of the SMEFT, involving a limited number of heavy fields, give rise to effective interactions aligned with the blind subspace defined in Eq. (II). We make use of MatchingDB [41] for this purpose.
@@
  • $\omega_{2}$, $\Omega_{2}$, $\mathcal{B}$ and $\mathcal{G}$ generate $\mathcal{O}_{dd}$.
@@
$\mathcal{O}_{dd}$$\{\omega_{2}\}\quad\{\Omega_{2}\}\quad\{\mathcal{G}\}$
@@
One might expect that finite one-loop matching effects, those not already captured by renormalisation group evolution, could lift the blind directions. However, we have verified explicitly that this is not the case for the single-scalar extensions; see Fig. 4. To this end, we performed one-loop matching onto the SMEFT using dedicated routines, with the assistance of SOLD [42, 43] and matchmakereft [44]. We note that integrating out massive vector fields at one loop is non-trivial without a complete description of the underlying spontaneously broken theory, and is currently not supported by existing automated tools [44, 45, 42, 43]. For brevity and cohesion, we omit the technical details of the matching procedure.
@@ IV Fit for multi-field models
The appearance of multiple leptoquarks is typical of GUTs and string-inspired models; see e.g. Ref. [46]. They are also common in models explaining flavour anomalies; see e.g. Refs. [47, 48, 49, 50].
@@
[…]
@@
Figure 6: Region of the parameter space excluded by EWPOs at $1\,\sigma$ including tree-level and leading-log RGEs (dark area enclosed by solid line) and including one-loop matching scalar corrections and RGE resummation (light area enclosed by dashed line) for a UV model involving two leptoquarks $\omega_{1}$ and $\mathcal{Q}_{1}$. The masses of the scalar and the vector are assumed equal and of $1$ TeV.
@@
Such a field content may be found in the left-right SUSY model, in which the SM gauge group is embedded into $SU(3)_{C}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$. The $W^{\prime}$ arises from the spontaneous breaking of $SU(2)_{R}$, while the second Higgs comes from the decomposition of Higgs bidoublets required by SUSY [51, 52].
@@
[…]
@@
[…]
@@
V Summary and Conclusions
@@
In this work, we have investigated blind directions in the SMEFT, focusing on the four-fermion sector with third-family WCs. By restricting our analysis to this sector, we have identified an 11-dimensional subspace that remains blind to EWPOs, even after accounting for improved perturbation theory. We have shown that simple multi-field UV completions of the SM, where the heavy fields are integrated out at tree level, can reside within this blind subspace. One-loop matching corrections do not alter the persistence of these blind directions.
@@
Our results make it clear that blind directions are not isolated curiosities but rather a systematic and generic feature of realistic SMEFT completions. Even under simplifying assumptions, such as limiting to third-generation couplings, working with tree-level matching, and considering only a modest number of operators, we identify multiple blind directions that evade the entire suite of EWPOs at TeraZ. The situation becomes even more pronounced when considering that 25 independent four-fermion operators (which can all be generated at tree level) are constrained by only 26 EWPOs, many of which are sensitive to overlapping subsets of operators. The apparent over-constrained nature of the system is therefore misleading. This imbalance, together with correlated SMEFT contributions from realistic multi-field UV completions, makes blind directions not just possible but likely. In this sense, our findings reveal that the problem is structural rather than exceptional, and underline the inherent limitations of relying solely on precision measurements, even at the unprecedented luminosities of TeraZ, to comprehensively probe new physics.
These results highlight the need for complementary high-energy collider experiments that can explore different kinematic regimes and help resolve the degeneracies present in precision electroweak measurements. This need becomes even more pressing as the typical scale of new physics increases beyond the 1 TeV benchmark used in our analysis. We also note that our fits are based on current experimental data rather than projected sensitivities for TeraZ. Nevertheless, our findings are in good agreement with previous studies where comparisons are possible. For example, despite differences in coupling assumptions, Fig. 3 of Ref. [3] shows that several single-field extensions remain unconstrained beyond 2 TeV.
It is also important to emphasize that the problem of blind directions becomes more severe when the simplifying assumptions used in this work are relaxed. Allowing for couplings to additional fermion families or permitting non-trivial hierarchies among WCs would further enlarge the space of blind directions, making it even more challenging to derive robust constraints from precision electroweak data alone. A comprehensive study of this more general scenario is left for future work.

### v2

@@ II Effective Field Theory Fit: bottom up
We emphasize that this is not a complete list of all blind directions, but just the simplest ones we have found with the stated assumptions. As we will see in the next sections, this is sufficient to illustrate potential blind spots at TeraZ, since there are several extensions of the SM that generate each of them.
We further observe that finite one-loop SMEFT effects [Biekotter:2025nln] slightly modify, but do not eliminate, the blind directions identified. In the following section, we systematically examine extensions of the SM that, when matched at tree level onto the SMEFT, give rise to effective operators aligned with these blind subspaces.
@@ III UV extensions of the SMEFT
The complete classification of SM extensions that generate dimension-six SMEFT operators at tree level was presented in Ref. [deBlas:2017xtg], building on earlier work [delAguila:2000rc, delAguila:2008pw, deBlas:2014mba]. For clarity, we reproduce in Table 1 only the subset relevant to our analysis, along with the corresponding effective operators they induce. We adopt the coupling notation used in Ref. [deBlas:2017xtg], and note that when multiple heavy fields are present, their contributions to the SMEFT Lagrangian simply add at leading order.
We assume that each heavy field $X$ has a single coupling to the SM, schematically of the form $X\psi_{1}\psi_{2}$, where $\psi_{1}$ and $\psi_{2}$ are SM fermions. The contribution of such a field $X$ to the SMEFT Lagrangian at tree level and dimension six is a linear combination of one or two Warsaw-basis operators with the field content $(\psi_{1})^{2}(\psi_{2})^{2}$. The only pairs of operators that may be generated are $\{\mathcal{O}_{x}^{(1)},\mathcal{O}_{x}^{(8)}\}$, with $x=qu$, qdqd and $ud$, and $\{\mathcal{O}_{lq}^{(1)},\mathcal{O}_{lq}^{(3)}\}$. The effects of the $\mathcal{O}_{x}^{(8)}$ operators on EWPO are all negligible. Thus, each heavy field we consider effectively generates a single operator relevant for EWPO, except for the ones that generate the $\mathcal{O}_{lq}$ operators. For the latter, we always consider heavy fields in pairs in which $c_{lq}^{(3)}$ is cancelled, so again only one operator (i.e. $\mathcal{O}_{lq}^{(1)}$) survives.
Our aim is to identify which tree-level completions of the SMEFT, involving a limited number of heavy fields, give rise to effective interactions aligned with the blind subspace defined in Eq. (II). We make use of MatchingDB [Aebischer:2023nnv] for this purpose.
@@
  • $\Omega_{2}$, $\mathcal{B}$ and $\mathcal{G}$ generate $\mathcal{O}_{dd}$.
@@
$\mathcal{O}_{dd}$ $\{\Omega_{2}\}\quad\{\mathcal{G}\}$
@@
One might expect that finite one-loop matching effects, those not already captured by renormalisation group evolution, could lift the blind directions. However, we have verified explicitly that this is not the case for the single-scalar extensions; see Fig. 4. To this end, we performed one-loop matching onto the SMEFT using dedicated routines, with the assistance of SOLD [Guedes:2023azv, Guedes:2024vuf] and matchmakereft [Carmona:2021xtq]. We note that integrating out massive vector fields at one loop is non-trivial without a complete description of the underlying spontaneously broken theory, and is currently not supported by existing automated tools [Carmona:2021xtq, Fuentes-Martin:2022jrf, Guedes:2023azv, Guedes:2024vuf]. For brevity and cohesion, we omit the technical details of the matching procedure.
@@ IV Fit for multi-field models
The appearance of multiple leptoquarks is typical of GUTs and string-inspired models; see e.g. Ref. [Croon:2019kpe]. They are also common in models explaining flavour anomalies; see e.g. Refs. [Assad:2017iib, DAmico:2017mtc, Calibbi:2017qbu, Cornella:2019hct].
@@
$\phantom{=}+y^{dq}_{\mathcal{Q}_{1}}\epsilon_{ABC}\mathcal{Q}_{1}^{A\mu\dagger}\bar{d}^{B}_{R}\gamma_{\mu}i\sigma_{2}q^{c\,C}_{L}+\text{h.c.}\,.$ (7)
@@
Figure 6: Region of the parameter space allowed by EWPOs at $1\,\sigma$ including tree-level and leading-log RGEs (dark area enclosed by solid line) and including one-loop matching scalar corrections and RGE resummation (light area enclosed by dashed line) for a UV model involving two leptoquarks $\omega_{1}$ and $\mathcal{Q}_{1}$. The masses of the scalar and the vector are assumed equal and of 11 TeV.
@@
Such a field content may be found in the left-right SUSY model, in which the SM gauge group is embedded into $SU(3)_{C}\times SU(2)_{L}\times SU(2)_{R}\times U(1)_{B-L}$. The $W^{\prime}$ arises from the spontaneous breaking of $SU(2)_{R}$, while the second Higgs comes from the decomposition of Higgs bidoublets required by SUSY [delAguila:2010mx, Hirsch:2015fvq].
@@
$\mathcal{L}_{UV}=y^{d}_{\varphi}\varphi^{\dagger}\bar{d}_{R}q_{L}+g^{du}_{\mathcal{B}_{1}}\mathcal{B}^{\dagger}_{1\,\mu}\bar{d}_{R}\gamma_{\mu}u_{R}+\text{h.c.}\,.$ (10)
@@
$\phantom{=}+g^{lq}_{\mathcal{U}_{2}}\mathcal{U}^{\mu\dagger}_{2}\bar{l}_{L}\gamma_{\mu}q_{L}+\frac{g_{\mathcal{X}}}{2}\mathcal{X}^{a\mu\dagger}\bar{l}_{L}\gamma_{\mu}\sigma^{a}q_{L}+\text{h.c.}\,.$ (13)
@@
V How things change at TeraZ
TeraZ will measure EWPD with far greater accuracy than LEP, and will therefore be sensitive not only to effects driven by the large top Yukawa but also to RGE contributions induced by gauge couplings.
As an example, consider the parameter space spanned by $c_{ud}^{(1)},c_{qd}^{(1)}$}. EWPD constrain these operators through their mixing into $c_{\phi b}$ and $c_{\phi q}^{(1)}$, which directly affect the observables $A_{b}$ and $A_{\text{FB}}^{b}$:
$\dot{c}_{\phi b}$ $\sim 6y_{t}^{2}(c_{qd}^{(1)}-c_{ud}^{(1)})+\frac{2}{3}g_{1}^{2}(c_{qd}^{(1)}+2c_{ud}^{(1)})\,$, (16)
$\dot{c}_{\phi q}^{(1)}$ $\sim-\frac{2}{3}g_{1}^{2}c_{qd}^{(1)}\,.$ (17)
At LEP precision, $g_{1}^{2}$ contributions can be neglected, which leads to the blind direction $c_{qd}^{(1)}\sim c_{ud}^{(1)}$. At TeraZ, however, $g_{1}^{2}$ effects are no longer negligible. Nevertheless, a blind direction persists, understood here as a region of parameter space where TeraZ sensitivity drops very significantly, and the same holds for most of the basic blind directions identified above; [^5] Note that, beyond these basic ones, exact blind directions will always exist, given that the parameter space has larger dimensionality than the number of EWPO. From an agnostic EFT standpoint, without additional knowledge about the UV, they cannot be simply dismissed as unlikely. see Fig. 9.
Figure 9: Fate of some of the LEP blind directions at TeraZ. Colour-code is as in Fig. 2. We have used the modified versions of smelli and flavio presented in Ref. [Allanach:2025wfi].
The only “exceptions” are the blind directions involving $c_{qe},c_{lq}^{(1)},c_{eu},c_{lu}$, not shown in the plots, for which couplings above $\sim 0.5$ can already be excluded at TeraZ. For all other cases, the blind directions merely narrow, and even couplings as large as $g\sim 0.5$ would remain beyond TeraZ sensitivity. This is also reflected in the fact that the simple SM extensions constructed earlier remain largely unconstrained by TeraZ data; see, for instance, Fig. 10.
Figure 10: TeraZ constraints at 22 $\sigma$ on the model of Fig. 8 using the same colour-coding.
One final comment is in order. Although our analysis assumes new physics at the TeV scale, running down to the electroweak scale via RGE, it is reasonable to expect that if such large couplings remain unconstrained, then new physics well below the TeV scale with $g\lesssim 1$ could also evade detection at TeraZ.
This limitation arises from working within EFT at a fixed scale, which is sensitive only to the ratio $g/\Lambda$ but not to $g$ and $\Lambda$ separately. This stands in contrast to hadron colliders. For instance, the particles in Fig. 10 could be searched for pair-produced dijet resonances [CMS:2024ldy] via gluon fusion, such as $gg\to\omega_{1}\omega_{1}$ with $\omega_{1}\to jj$. The small coupling to quarks, undetectable at TeraZ, would merely increase the $\omega_{1}$ lifetime, without otherwise diminishing the sensitivity of hadron collider searches. This further motivates later high-energy stages of the FCC programme, including $e^{+}e^{-}$ runs at different energies.
VI Summary and Conclusions
@@
In this work, we have investigated blind directions in the SMEFT, focusing on the four-fermion sector with third-family WCs. By restricting our analysis to this sector, we have identified high-dimensional subspace that remains blind to EWPOs, even after accounting for improved perturbation theory. We have shown that simple multi-field UV completions of the SM, where the heavy fields are integrated out at tree level, can reside within this blind subspace. One-loop matching corrections do not alter the persistence of these blind directions.
@@
Our results make it clear that blind directions are not isolated curiosities but rather a systematic and generic feature of realistic SMEFT completions. Even under simplifying assumptions, such as limiting to third-generation couplings, working with tree-level matching, and considering only a modest number of operators, we identify multiple blind directions that evade the entire suite of EWPOs at TeraZ. The situation becomes even more pronounced when considering that 25 independent four-fermion operators (which can all be generated at tree level) are constrained by only 26 EWPOs, many of which are sensitive to overlapping subsets of operators. Our findings reveal that the problem is structural rather than incidental, highlighting limitations of relying solely on precision measurements, even at the unprecedented luminosities of TeraZ, to comprehensively probe new physics.
These results provide further motivation for complementary high-energy collider experiments. Future FCC-ee runs at higher center-of-mass energies and the FCC-hh will open access to new kinematic regimes and help resolve the degeneracies that remain in precision electroweak measurements. Their role becomes especially valuable as the typical scale of new physics might extend beyond the 1 TeV benchmark considered in our analysis.